Each of the four tires of Coach Carsons ride is filled with a different gas, to a tire gauge of 3.0 atm at 83 F The mass o
Answer #1

Answer - Given , Mass of Tire 1 = 116 g (Air)

Tire 2= 80.7 g(Neon)

Tire 3 =16 g (Helium)

Tire 4= 160 g (unknown)

Pressure =3.0 atm

Temperature = 83 F

for tire 2

the no. of molecules/atoms = (80.7/20.18)*6.02 *1023 =24.07 * 1023 atoms/ molecules

no. of moles = 3.99 moles

also PV=nRT

3 *V = 3.99 * 0.08206 * 301.48

V = 32.90 L it will be same for all tires

Now,

(a) We know that.

PV=nRT

where, P = Pressure

V = Volume (32.90L calculated above)

n = no. of moles

R = gas constant = 0.08206 L atm K-1mol-1

T = temperature (in kelvin) =301.483

3*32.90 = n *0.08206* 301.483

n =3.989 moles

So no. of atoms in 0.124 moles = 3.989 * 6.02 * 1023 = 24.02 * 1023 atoms

the no. of atoms/ molecules is 24.02 *1023 atoms/ molecules and is same for all the tires.

(b) for tire 4 we have to find the molecular mass of the unknown gas

The no. of atoms/ molecules is same i.e. 24.02 *1023 atoms/ molecules

  

We know that,

No. of atoms/ molecules = (mass/ molecular mass) * 6.02 *1023

Put all the values

24.02* 1023 = (160/ molecular mass) * 6.02 * 1023

molecular mass = (160 * 6.02) / 24.02

= 40.09 u

So, the molecular mass of unkown gas is 40.09u

(c) The molecular mass of Argon is 39.948u which almost equal to our answer 40.09u. So, the unknown gas in the Tire 4 must be ARGON.